# Hierarchical Reasoning Model: A Critical Supplementary Material

### v1

@@ Abstract
Transformers have demonstrated remarkable performance in natural language processing and related domains, as they largely focus on sequential, autoregressive next-token prediction tasks. Yet, they struggle in logical reasoning, not necessarily because of a fundamental limitation of these models, but possibly due to the lack of exploration of more creative uses, such as latent space and recurrent reasoning. An emerging exploration in this direction is the Hierarchical Reasoning Model Wang et al. (2025a), which introduces a novel type of recurrent reasoning in the latent space of transformers, achieving remarkable performance on a wide range of 2D reasoning tasks. Despite the promising results, this line of models is still at an early stage and calls for in-depth investigation. In this work, we perform a critical review on this class of models, examine key design choices and present intriguing variants that achieve significantly better performance on the Sudoku-Extreme and Maze-Hard tasks than previously reported. Our results also raise surprising observations and intriguing directions for further research.
@@ 2 Latent Space Reasoning and Hierarchical Reasoning Model (HRM)
The Hierarchical Reasoning Model (HRM) (Wang et al., 2025a) is the latest popular model that performs latent-space reasoning. It introduces a dual-loop recurrent architecture inspired by cognitive science. It alternates between a fast low-level (L) module that executes candidate reasoning steps and a slower high-level (H) module that updates a more abstract plan. HRM iterates these cycles until a learned halting signal indicates completion, yielding effective depth far beyond its nominal layer count. Notably, HRM achieves strong performance on reasoning-heavy benchmarks such as Sudoku and ARC-AGI-1 (Chollet, 2019) without pretraining or explicit chain-of-thought supervision (Wang et al., 2025a). Despite the promising ideas and concepts introduced by the paper, it is unclear how much benefit these L and H modules bring, compared to a plain transformer. In our experiments in Figure 2(b), we show that even without H module, a plain 8-layer L-Module only HRM performs similarly to the original 4-layer L-module, 4-layer H-module HRM. This is consistent to what is reported by ARC-Foundation (2025), which says that the plain transformer performs similarly to HRM.
@@ 3.1 HRM is a Diffusion Model
4 Adaptive Recurrent Reasoning for Both Training and Inference
4.1 Key Finding
We identify a critical mismatch in the original HRM implementation: while its Adaptive Computation Time (ACT) module learns to decide when to halt during training, evaluation ignores this policy and always executes the maximum number of reasoning steps. To address this, we modify HRM to use its learned halting policy during evaluation as well.
Our adjustment allows HRM at evaluation time to more faithfully reflect the model’s learned reasoning process. By accessing the learned ACT during inference, the model is able to decide earlier on that a puzzle is completed, preventing potential errors it could make if it were to complete its maximum number of reasoning steps. Indeed, as we observe in Figure 1 and Figure 2, our modified HRMs highly benefit from applying ACT halting during inference, greatly surpassing the original HRM performance. The similarity in performance between Figure 1, where we allow the 4-layer H- and L-module HRM to use ACT halting during evaluation, and Figure 2, where we allow an 8-layer L-module HRM to use ACT halting during evaluation, casts doubt on whether the biologically-inspired hierarchical H- and L- module design choice provides much benefit. In Figure 3, we observe that the model flexibly allocates computation steps across different puzzles, with some puzzles being solved in very few reasoning steps. Our findings also raise interesting questions about how models can reason, and how they differ from or align with human reasoning.
@@ 4.2 Methods
We build on the Hierarchical Reasoning Model (HRM) introduced by Wang et al. (2025a), which integrates the ACT module into a multi-step recurrent reasoning architecture. The HRM architecture consists of two hierarchical reasoning modules: an H module that processes global abstractions, updating less frequently, and an L module that handles local reasoning, which updates more frequently. The model also maintains an inner carry state that preserves computational context across reasoning iterations at every step. The model’s ACT module uses Q-learning to balance exploration (continuing reasoning) against exploitation (terminating with the current solution). At each step , a Q-head outputs values for halt and continue as . Halting occurs if (i) the maximum horizon is reached or (ii) once a stochastic minimum threshold has been satisfied.
@@
We modify HRM to apply halting learned during training to evaluation as well. Specifically, a sequence is completed when it either halts naturally or reaches the maximum step limit . We use rather than the stochastic comparison in the original HRM training implementation to determine whether to halt or continue the sequence, as evaluation does not benefit from using the exploratory logits.
@@
Our changes ensure that evaluation tests the learned halting policy, not just maximum-depth reasoning of the model. We compare our HRM, where ACT is used in both training and evaluation, against the original HRM with fixed-step evaluation to measure accuracy improvements. Specifically, we train and evaluate this modified HRM, which utilizes its halting capabilities in the evaluation mode, on the released Sudoku-Extreme and Maze-Hard datasets.
@@ 4.3 Results
Remarkably, our modified HRM implementation, which uses ACT in both training and evaluation, largely outperforms the original HRM, achieving nearly perfect performance on the Sudoku-Extreme dataset compared to the original HRM’s performance (Figure 1). This shows that halting during evaluation is a crucial component for realizing the model’s reasoning potential. When the model continues reasoning after it has already produced a correct solution, additional steps can introduce errors. Over-reasoning can harm performance, as the model’s fixed-step evaluation can force it to keep making edits and predictions even when those updates are unnecessary. By contrast, allowing the ACT policy to halt during inference can help prevent this drift. We can observe more of the model’s flexible halting process, and its ability to reason in fewer steps during evaluation in Figure 3.
In Figure 2, we remove the H-module in the HRM and instead expand the L-module to 8 layers, while continuing to enable ACT during evaluation. This model, with the ability to use ACT during evaluation, also consistently achieves higher accuracy than the original HRM, while also requiring fewer steps on average to complete a puzzle. Despite this modified HRM using only a single 8-layer L-module, doing away with biologically-inspired hierarchical design choice of the fast-updating L-module and slow-updating H-module, our achieves similar performance as our experiments with the original L- and H-module HRM in Figure 1. From these similar performance results, it is unclear how much benefit the hierarchical L and H modules bring to the HRM.
4.4 Reasoning with Few Steps
Figure 3 provides concrete examples where the model halts after only a handful of reasoning steps, while still producing correct solutions. The model dynamically adjusts computation across puzzles and sometimes solves difficult puzzles with very short trajectories. This behavior contrasts with human reasoning, where solving a Sudoku puzzle typically involves many step-by-step constraint propagations until the grid is fully filled. Adaptive halting allows the model not only to avoid unnecessary computation but also to achieve correct solutions through compressed reasoning pathways, raising interesting questions about how neural models represent and execute reasoning.
5 Generalization to Other Datasets
5.1 Maze-Hard
To show that our approach generalizes to other datasets, we tested its performance on the Maze-Hard dataset. The results are shown in Figure 4. Our approach converges much faster and doubles the HRM’s performance within 30,000 training iterations.
5.2 ARC-AGI
ARC-AGI has a large board size and requires significantly more computational resources to run, which can take a few days on a single GPU. At the time of writing, we are actively running experiments on ARC-AGI. We will report results on ARC-AGI in a future revision of the paper.
6 Discussion
Our work highlights an interesting and strong phenomenon that as the model learns better, it solves the problem with fewer steps. It stands in stark contrast to the common belief that more reasoning steps during inference would lead to better performance. As the HRM paper repeatedly states in the “Inference-time scaling” section: “An effective neural model should exploit additional computational resources during inference to enhance performance … Models trained with a specific can generalize to higher computational limits during inference, leading to improved accuracy” (Wang et al., 2025a). As what we see in our results, if given the chance, a well trained model actually choose to solve problems with less steps, which is very counterintuitive and intriguing.
As another related investigation, ARC-Foundation (2025) published a blog post discussing many aspects about HRM, pointing out that the ACT plays an important factor in HRM’s good performance. However, ARC-Foundation (2025) does not observe much performance differences between any settings with ACT loop 4 or even 2. They also do not observe the importance of ACT during inference, nor any out-of-curve performances as we see. In fact, ARC-Foundation (2025) states: “This indicates that training with refinement loops is more important than using the refinement for inference.” This conclusion is exactly the opposite of what we show in this paper, where inference time ACT makes a significant difference.
@@ 7 Conclusion
In this paper, we perform a critical review on the recently popular reasoning model HRM. We show that HRM’s architecture can be largely simplified to a 8-layer plain transformer while maintaining good performance. We demystify HRM’s “one-step approximation of BPTT” by showing its equivalence to diffusion/consistency models. We explore variants of HRM and achieve drastically better performance than the original HRM model, raising the Sudoku-Extreme performance from previously reported 55% to above 90%. We highlight the interesting relationship between testing performance and testing reasoning steps, where the model learns to reason with fewer and fewer steps while achieving increasingly better performance. Visualization on Sudoku examples provides an intuitive demonstration of model’s surprising behavior, which differs greatly from human problem solving.
Acknowledgments
This work was supported by the K. Lisa Yang Integrative Computational Neuroscience (ICoN) Center, one of the six centers of the Yang Tan Collective at MIT. This work was also supported by the Center for Brains, Minds and Machines (CBMM), funded by NSF STC award CCF - 1231216.

### v2

# Hierarchical Reasoning Models: Perspectives and Misconceptions

@@ Abstract
Transformers have demonstrated remarkable performance in natural language processing and related domains, as they largely focus on sequential, autoregressive next-token prediction tasks. Yet, they struggle in logical reasoning, not necessarily because of a fundamental limitation of these models, but possibly due to the lack of exploration of more creative uses, such as latent space and recurrent reasoning. An emerging exploration in this direction is the Hierarchical Reasoning Model Wang et al. (2025a), which introduces a novel type of recurrent reasoning in the latent space of transformers, achieving remarkable performance on a wide range of 2D reasoning tasks. Despite the promising results, this line of models is still at an early stage and calls for in-depth investigation. In this work, we review this class of models, examine key design choices, test alternative variants and clarify common misconceptions.
@@ 1 Introduction
Despite the significant attention gained by HRM due to its innovative designs and impressive performance on various reasoning benchmarks, many aspects remain unclear. This paper aims to investigate some key design elements of HRM, for example: 1. Is HRM’s hierarchical architecture important, in particular the L and H modules? 2. How novel is HRM’s “biologically plausible” one-step gradient training? 3. What benefit does HRM’s Adaptive Computation Time (ACT) bring? Why is there a discrepancy in HRM’s training and test procedures? Through analyses of these key features, we aim to clarify common misconceptions regarding their importance and offer a clearer understanding of the model.
@@ 2 Latent Space Reasoning and Hierarchical Reasoning Model (HRM)
The Hierarchical Reasoning Model (HRM) (Wang et al., 2025a) is the latest popular model that performs latent-space reasoning. It introduces a dual-loop recurrent architecture inspired by cognitive science. It alternates between a fast low-level (L) module that executes candidate reasoning steps and a slower high-level (H) module that updates a more abstract plan. HRM iterates these cycles until a learned halting signal indicates completion, yielding effective depth far beyond its nominal layer count. Notably, HRM achieves strong performance on reasoning-heavy benchmarks such as Sudoku and ARC-AGI-1 (Chollet, 2019) without pretraining or explicit chain-of-thought supervision (Wang et al., 2025a). Despite the promising ideas and concepts introduced by the paper, it is unclear how much benefit these L and H modules bring, compared to a plain transformer. In our experiments in Figure 1, we show that even without H module, a plain 8-layer L-Module only HRM performs similarly to the original 4-layer L-module, 4-layer H-module HRM. This is consistent to what is reported by ARC-Foundation (2025), which says that the plain transformer performs similarly to HRM.
@@ 3.1 HRM is a Diffusion Model
4 Adaptive Computation Time (ACT) in HRM
Adaptive Computation Time (ACT), in a general sense, is a concept of allowing a model to perform different amount of computation (e.g., number of iterations) on different input data. It was proposed in (Graves, 2016) on Recurrent Neural Networks (RNN). For a feedforward network, applying some form of ACT essentially converts it into an RNN and allows it to scale its computation beyond a feedforward model and potentially achieve Turing Completeness (Liao & Poggio, 2016). This is one of the core motivations of HRM, which can be interpreted as roughly equivalent to applying ACT to transformers.
Although ACT is a key design in HRM, it is not sufficiently discussed in the original paper (Wang et al., 2025a). There is a critical mismatch in the original HRM implementation: while its ACT module learns to decide when to halt during training, evaluation ignores this policy and always executes the maximum number of reasoning steps. To further explore this, we modify HRM to use its learned halting policy during evaluation as well.
There are a few different ways of performing ACT during inference: 1. baseline: apply a fixed but different number of reasoning steps from training for all examples 2. use a decision rule to decide the number of reasoning steps for every sample. We will describe them in detail in the experiments section.
5 Experiments
@@ 5.1 Methods
5.1.1 Removal of H-Module
The original HRM alternates between two reasoning modules: a fast-updating low-level L-module and a slower high-level H module that updates once every steps. These modules maintain internal latent states and are updated as:
In Wang et al. (2025a), the intuition of these hierarchical reasoning modules is that the H-module updates less frequently, as it is meant to process global abstractions, while the L-module updates more frequently, as it is intended to handle local reasoning. However, in our experiments, we create an expanded 8-layer L-module only HRM that updates at each timestep , eliminating the biologically-inspired hierarchical design choice of maintaining two separate latent states and . As shown in Figure 1, our approach achieves similar performance as the original HRM, casting doubt on how much benefit the hierarchical L and H modules bring to the HRM.
5.1.2 ACT for both Training and Inference
We also investigate the HRM’s ACT module. The model’s ACT module uses Q-learning to balance exploration (continuing reasoning) against exploitation (terminating with the current solution). At each step , a Q-head outputs values for halt and continue as . Halting occurs if (i) the maximum horizon is reached or (ii) once a stochastic minimum threshold has been satisfied.
@@ 5.1.2 ACT for both Training and Inference
We modify HRM to apply halting learned during training to evaluation as well, determined by the halting logits ( and ). We experiment (as shown in Figure 3) with halting the model in two cases: (1) when , and (2) when , observing the effects of varying the halting threshold on model performance.
@@ 5.2 Results
In Figure 1, we remove the H-module in the HRM and instead expand the L-module to 8 layers.
Figure 5 provides concrete examples where the model halts after only a handful of reasoning steps, while still producing correct solutions. The model dynamically adjusts computation across puzzles and sometimes solves difficult puzzles with very short trajectories. This behavior contrasts with human reasoning, where solving a Sudoku puzzle typically involves many step-by-step constraint propagations until the grid is fully filled. This raises interesting questions about how neural models represent and execute reasoning.
@@ 6 Conclusion
In conclusion, our investigation reveals several critical insights into the HRM model. First, we found that the H module does not significantly contribute to overall performance; in fact, an HRM model utilizing only the L module performs comparably to the original HRM, consistent to what is reported by ARC-Foundation (2025). Second, the one-step gradient approximation employed by HRM is effectively equivalent to the training paradigm used in diffusion models. Finally, our analysis indicates that ACT does not enhance inference performance; rather, running the model for the maximum number of reasoning steps yields the best results. This raises an intriguing question regarding the recurrent nature of HRM: is it genuinely recurrent, or does it function similarly to a very deep feedforward model network. These findings challenge existing assumptions about HRM and suggest avenues for further research.